OpenStudy (crashonce):

The interior angles of a convex polygon are in AP. If the largest angle is 140, and the smallest is 100, then the number of sides of the polygon is? 6, 8, 10, 11 or 12

OpenStudy (crashonce):

@ikram002p

OpenStudy (crashonce):

@ganeshie8

OpenStudy (crashonce):

@jim_thompson5910

OpenStudy (crashonce):

@Compassionate

ganeshie8 (ganeshie8):

\[100 + (n-1)d = 140\] \[(n-1)d = 40\]

OpenStudy (crashonce):

not sure

OpenStudy (ikram002p):

i might confsed u xD i remove my comments and read carfully what ganesh wrote

OpenStudy (crashonce):

so its 11?

ganeshie8 (ganeshie8):

Also the interior angles must add up to : \(180(n-2)\)

ganeshie8 (ganeshie8):

So you have two equations basically : \[(n-1)d = 40\tag{1}\] \[\frac{n}{2}\left[2\times 100 + (n-1)d\right] = 180(n-2)\tag{2}\]

OpenStudy (crashonce):

wait is it 6 or 8

ganeshie8 (ganeshie8):

you have two equations and two unknowns, you can solve them right ?

OpenStudy (crashonce):

ooooo its 6 right? sorry

ganeshie8 (ganeshie8):

yes let me knw if something doesnt make sense..

OpenStudy (crashonce):

it makes sense i gott another one

ganeshie8 (ganeshie8):

we have used "nth term" and the "partial sum" formulas of AP